1- How many ways are there to select a committee of 50% of 16 from 80% of the 110 people and select 50% of 16 for separate positions from 20% of 110 people

please with all the steps.
Answer #1

50% of 16=8 80% of 110=88 20% of 110=22

so if you simplify the question you will get:

How many ways are there to select a committee of 8 from 88 people and select 8 for separate positions from 22 people.

Now, the number of ways of selecting 8 people from 88= and the number of ways of selecting 8 people from 22=

so total number of ways=
Answer #2

Problem Breakdown:

  1. First Group (80% of 110 people):

    • Number of people = 0.8×110=88.

    • Committee size = 0.5×16=8.

    • Ways to choose:

      (888)

  2. Second Group (20% of 110 people):

    • Number of people = 0.2×110=22.

    • Committee size = 0.5×16=8.

    • Ways to choose:

      (228)

  3. Total Ways:
    Multiply the two combinations (since selections are independent):

    Total=(888)×(228)

Final Answer:

Note:

  • (nk) = "n choose k" = n!k!(nk)!.

  • No simplification needed unless numerical values are required.


answered by: anonymous